OpenStudy (anonymous):

What is the solution to the equation 4^(2x) ≈ 3 ? x = –0.631 x = –0.396 x = 0.396 x = 0.631

OpenStudy (anonymous):

take logs of both sides.

OpenStudy (anonymous):

Also, use the relationship:\[\ln a ^{b} = b \times \ln a\]

OpenStudy (anonymous):

no need of logs just replace x with the value given and solve

OpenStudy (anonymous):

@Sufiyancs , that's no way to get an answer! What if you were not given any choices! The goal here is not to get the answer, it's to learn how to solve problems!

OpenStudy (anonymous):

fine u continue man.. i'm out

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[4^{2x} = 3\]can be solved by\[\ln 4^{2x} = \ln 3\]which equals\[2x \times \ln 4 = \ln 3\]

OpenStudy (anonymous):

Divide each side by 2(ln 4) and you get your value for "x"

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

Can you get the value for ln 3 ?

OpenStudy (anonymous):

what do you mean by ln 3?

OpenStudy (anonymous):

Just what everyone in the world of math means by it. ln 3 is simply ln 3, the natural logarithm of 3. The exponent you put on "e" to get 3. Logs are exponents.

OpenStudy (anonymous):

Do you have any familiarity with logarithms?